simplify the following expression
(3 - 2√7)(3 + 2√7)

Answer by texttutoring(324)

Do FOIL.
(3-2sqrt7)(3+2sqrt7)
=9 -4(sqrt7)(sqrt7))
=9-4(7)
=9-28
=-19